How many cubic cm of iron are there in an open box whose external dimensions are 36cm, 25 cm and 16.5 cm. the iron being 1.5 cm thick throughout. If 1 cubic cm of iron weighs 15g, find the weight of the empty box in Kgs

Grade:

1 Answers

Nishant Vora IIT Patna
9 years ago
Volume of cube is = 36*25*16.5 = 14850 cm3.
1 cc = 15 grams
so 14850 cc = ? grams

Now use unity method
weight of iron = 14850*0.015Kg = 222.75 Kg

I hope its clear